# Adversarial Inception for Bounded Backdoor Poisoning in Deep Reinforcement Learning

### v1

@@ Appendix A Appendix
[…]
2Contents
@@ A.1 Proofs for Adversarial Inception Theoretical Guarantees
In this section we provide proofs of all the results we claim in Section 4 – first proving Theorem 1, proceeding through proofs of Lemmas 0 through 2, and subsequently proving Theorem 2. It should be noted that Lemma 0 is not a main claim of our paper and is instead a convenient result adapted from part of a proof by Rathbun et al. (2024).
@@ A.1.2 Lemma 0
This is labeled Lemma 0 as it is a useful result which will be used in both Lemma 1 and Lemma 2, but isn’t a key result for this paper. It should be noted that the derivation is identical to one seen in Rathbun et al. (2024), though here we are generalizing and replicating the result so it can be referenced with confidence in Lemma 1 and Lemma 2
@@ A.1.3 Lemma 1
From here, similar to Rathbun et al. (2024), we note that the second summation is over $s^{\prime}\in S_{p}$, yet the term is always inverted with $\delta^{-1}$. Since $\delta$ is bijective we can therefore convert the summation to one over $s^{\prime}\in S$:
@@ A.1.4 Lemma 2
Due to the nature of this proof, many of the steps are nearly identical to the proof given for Lemma 1, with some minor notational differences (using $\pi$ instead of $\pi^{*}$). Thus we will provide an abridged version of the proof, with citations to the relevant steps from Lemma 1 when relevant. Thus we quickly derive an intermediate result similar to 37:
@@ A.2 More Experimental Details and Hyper Parameters
In this section we give further details on the hyper parameters and setups we used for our experimental results. First, in Table 6 we summarize each environment we studied, their properties, and the learning parameters we used in each experiment. Parameters not mentioned in the table are simply default values chosen in the cleanrl (Huang et al., 2022) implementation of PPO.
@@
Q*Bert Video Game Image 15M 0.00025 QbertNoFrameskip-v4
@@
Table 6: Further details for each environment tested in this work. All action spaces were discrete in some form, though for Safety Car a discretized versions of its continuous action space was used. The “Environment Id.” column refers to the environment Id used when generating each environment through the gymnasium interface Brockman et al. (2016).
Next in Table 7 we provide the generic attack parameters used for each attack in our experiments. Across all our image based domains we utilized a 6x6, checkerboard pattern of 1s and 0s in the top left corner of the image as a trigger. For Safety Car we set 4 values of the agent’s lidar sensors, corresponding to their relative cardinal directions, to be equal to 1 indicating 4 objects placed directly in front of, behind, to the left, and to the right of the agent. This type of pattern is not possible in this environment as only one of the associated object, dubbed “vases”, exist in SafetyCarGoal1. Lastly, for CAGE-2 we simply append a boolean value to the end of the agent’s observation which we set to 1 in poisoned states, or 0 otherwise. This environment in particular has a very simple observation space, being a 52 bit, one-hot encoding of the environment. Due to this, it’s unclear how one would best devise a trigger pattern that meets our assumption of $S$ and $S_{p}$ being disjoint. Therefore we chose a to use a trigger which cannot be set to 1 in any other case than a poisoned state. Abstractly this can be seen as representing some malicious or otherwise strange network behavior which might be represented in the observations of an agent trained on a real-life corporate network.
Values for $\beta_{low}$ and $\beta_{high}$ were chosen to balance attack success and episodic return. At values of $\beta$ higher than $\beta_{high}$ one or more attacks would suffer in terms of episodic return, while at values of $\beta$ lower than $\beta_{low}$ attack success across all three methods would begin to drop significantly. $\beta$ values were chosen on a per-environment basis using the parameters chosen by Rathbun et al. (2024) as a starting point.
Attack Details
Environment Trigger $\beta_{low}$ $\beta_{high}$ Target Action
Q*bert Checkerboard Pattern 0.01% 0.03% Move Right
Highway Merge Checkerboard Pattern 7.5% 10% Merge Right
Safety Car Lidar Pattern 0.05% 0.1% Accelerate
CAGE-2 Boolean Indicator 0.5% 1.0% Sleep (No-Op)
Table 7: Attack and learning parameters used for each environment. $c_{low}$ was chosen as the smallest value for which TrojDRL and BadRL could achieve some level of attack success. $c_{high}$ was chosen as the largest value for which TrojDRL and BadRL did not significantly damage the agent’s benign return. A similar method was used in determining the poisoning budget.
@@ A.2.2 Q-Incept Attack Parameters
Environment Steps per Update Start Poisoning Threshold
Q*bert 50 6.7%
Highway Merge 2 10%
Safety Car 4 4.0%
CAGE-2 4 4.0%
@@
A.3 Further Experimental Results and Analysis
A.3.1 Numerical Results over Different Poisoning Budgets
Empirical Comparison of Backdoor Attacks Under Bounded Rewards
Environment CAGE-2 ($\beta=0.5\%$) Highway Merge ($\beta=7.5\%$) Qbert ($\beta=0.01\%$) Safety Car ($\beta=0.1\%$)
Mertric ASR $\sigma$ ASR $\sigma$ ASR $\sigma$ ASR $\sigma$
Q-Incept 30.61% 15.01% 53.03% 25.56% 100.00% 0.00% 100.00% 0.00%
SleeperNets 0.00% 0.00% 1.47% 0.84% 19.98% 4.39% 83.95% 13.45%
TrojDRL 0.00% 0.00% 3.27% 3.56% 15.38% 3.05% 53.35% 9.51%
Mertric BRR $\sigma$ BRR $\sigma$ BRR $\sigma$ BRR $\sigma$
Q-Incept 92% 17.71% 98.27% 0.84% 100% 8.07% 76.51% 7.35%
SleeperNets 72.00% 18.55% 99.92% 0.07% 100% 8.66% 91.42% 5.97%
TrojDRL 100% 21.25% 99.86% 0.13% 100% 13.24% 86.37% 9.65%
Table 9: Numerical results for each attack under lower poisoning rates $\beta_{low}$.
In Table 9 we provide further numerical results for the performance of each attack under lower poisoning rates than those presented in the main body. In the cases of CAGE-2, Highway Merge, and Safety Car we can see that the performance of SleeperNets and TrojDRL did not meaningylfully change. This indicates that these attacks likely will not improve given poisoning rates even higher than $\beta_{high}$. In the case of Q*Bert, the ASRs of SleeperNets and TrojDRL do drop off a decent amount from the original poisoning rate of $0.03\%$ presented in the main body, so it is possible they can find better performance as $\beta$ scales.
Q-Incept was fairly consistent in ASR between $\beta_{low}$ and $\beta_{high}$ on all environments excluding CAGE-2. This large drop in ASR indicates that a poisoning rate of $0.5\%$ is too low for Q-Incept to be successful under bounded reward poisoning on CAGE-2. Across all environments and between $\beta_{low}$ and $\beta_{high}$ the BRR performance of each attack did not meaningfully change.
A.3.2 Training Plots for Q*Bert and Safety Car
Here in Figure 6 and Figure 8 we present the training curves for TrojDRL, Q-Incept, and SleeperNets on the Safety Car and Q*Bert environments respectively. We can see that all attacks perform similarly over time in terms of episodic return on both environments, but Q-Incept is the only attack to reach 100% ASR on average in both environments – doing so very quickly.
Figure 6: Performance of TrojDRL, Q-Incept, and SleeperNets on the Safety Car environment in terms of ASR and episodic return.
Figure 7: Performance of TrojDRL, Q-Incept, and SleeperNets on Q*Bert in terms of ASR and episodic return.
A.3.3 How Often Does Q-Incept Alter the Agent’s Actions?
Here we explore how often the Q-Incept attack alters chosen actions in the agent’s replay memory $\mathcal{D}$. In Figure 8 we see that the attack generally balances its action poisoning over time, altering actions on roughly 50% of the time steps it poisons. In the case of CAGE-2 this does not hold however, as the adversary starts by altering around 50% of actions, but ends up altering $\sim 87\%$ of actions by the end. To us this indicates that the difference in values between good and bad actions was much larger in CAGE-2 than in other environments, and furthermore that the agent was highly likely to choose these actions over others as training progressed. Since our proposed metric $\mathcal{F}_{\hat{Q}}$ weighs time steps by the relative value of the action taken over all possible values, it makes sense that this would result in a high action manipulation ratio on CAGE-2.
Figure 8: Ratio of actions changed across poisoned states in all four environments. Values are measured as the number of actions changed divided by the total number of timesteps the attack has poisoned.
A.4 Further Discussion
@@ A.4.1 Motivation for Baselines
The only other, somewhat comparable attack is BadRL (Cui et al., 2023) which builds upon TrojDRL by optimizing the adversary’s trigger pattern to achieve greater attack success. Trigger optimization is effective but orthogonal to the goals of this work as it can be generically applied to any attack. Furthermore, Rathbun et al. (2024) showed that BadRL without this trigger fine-tuning often performs worse than TrojDRL, likely since it uses methods to poison the most important – and thus hardest to poison – states in the MDP. Taking all of this into consideration we decided to omit BadRL from our empirical study. Therefore TrojDRL is the best baseline to use when comparing against static reward poisoning attacks using forced action manipulation.
A.4.2 Motivation for Environments
In this paper we study 4 environments – Q*Bert, Safety Car, CAGE-2, and Highway Merge. In the TrojDRL paper the authors focused their empirical studies towards Atari game tasks in the gym API Brockman et al. (2016). We think it is useful to include some of these environments like Q*Bert, as they are standard baselines for RL in general, however we believe it is critical to extend this study to further domains when studying the potential impacts of adversarial attacks. This belief is supported by the findings of Rathbun et al. (2024) who showed that TrojDRL – which consistently attains near 100% ASR on Atari environments without bounded reward poisoning constraints – often fails to achieve high ASR when tested on non-Atari environments.

### v3

@@ Appendix A Appendix
Section Contents
Appendix A.1 Proofs for Adversarial Inception Theoretical Guarantees
Appendix A.1.1 Theorem 1
Appendix A.1.2 Lemma 0
Appendix A.1.3 Lemma 1
Appendix A.1.4 Lemma 2
Appendix A.1.5 Theorem 2
Appendix A.2 More Experimental Details and Hyper Parameters
Appendix A.2.1 TrojDRL and SleeperNets Parameters
Appendix A.2.2 Q-Incept Attack Parameters
Appendix A.3 Further Discussion
Appendix A.3.1 Motivation for Baselines
Appendix A.3.2 Motivation for Environments
Appendix A.3.3 Comments on Compute Overhead
Appendix A.4 Further Experimental Results and Analysis
Appendix A.4.1 Q-Incept Against DQN
Appendix A.4.2 Q-Incept at Different Poisoning Rates
Appendix A.4.3 Further Ablations
Appendix A.4.4 Training Plots for Remaining Environments
Appendix A.4.5 How often does Q-Incept Alter the Agent’s Actions?
@@ A.1 Proofs for Adversarial Inception Theoretical Guarantees
Attack Q-Incept (Ours) SleeperNets TrojDRL BadRL
Reward Poisoning Constrained Dynamic Static Static
Action Manipulation Inception None Forced Forced
Attack Success Guarantees of Existing Backdoor Attacks in DRL
Unconstrained Rewards Yes Yes No No
Constrained Rewards Yes No No No
Table 4: Summary of the currently known attack success guarantees of different backdoor attack strategies. New results proven in this paper are printed in bold. Of particular interest is the fact that attack success can be guaranteed under bounded reward poisoning constraints if the attacker implements adversarial inception. Without inception there are no guarantees. Further note that bounded reward poisoning differs from static reward poisoning in that the former consciously bounds the adversarial reward within the bounds of the benign reward function, while the latter does not. Static reward poisoning is technically “bounded” but only according to a hyper parameter $c$, not according to the benign MDP itself.
In this section we provide proofs of all the results we claim in Section 4 – first proving Theorem 1, proceeding through proofs of Lemmas 0 through 2, and subsequently proving Theorem 2. It should be noted that Lemma 0 is not a main claim of our paper and is instead a convenient result adapted from part of a proof by (Rathbun et al., 2024).
@@ A.1.2 Lemma 0
This is labeled Lemma 0 as it is a useful result which will be used in both Lemma 1 and Lemma 2, but isn’t a key result for this paper. It should be noted that the derivation is identical to one seen in (Rathbun et al., 2024), though here we are generalizing and replicating the result so it can be referenced with confidence in Lemma 1 and Lemma 2
@@ A.1.3 Lemma 1
From here, similar to (Rathbun et al., 2024), we note that the second summation is over $s^{\prime}\in S_{p}$, yet the term is always inverted with $\delta^{-1}$. Since $\delta$ is bijective we can therefore convert the summation to one over $s^{\prime}\in S$:
@@ A.1.4 Lemma 2
Due to the nature of this proof, many of the steps are nearly identical to the proof given for Lemma 1, with some minor notational differences (using $\pi$ instead of $\pi^{*}$). Thus we will provide an abridged version of the proof, with citations to the relevant steps from Lemma 1 when relevant. Thus we quickly derive an intermediate result similar to 36:
@@ A.2 More Experimental Details and Hyper Parameters
In this section we give further details on the hyper parameters and setups we used for our experimental results. In Table 5 we summarize each environment we studied, their properties, and the learning parameters we used in each experiment. Parameters not mentioned in the table are simply default values chosen in the cleanrl (Huang et al., 2022) implementation of PPO.
@@
Q*Bert Atari Game Image 15M 0.00025 QbertNoFrameskip-v4
Frogger Atari Game Image 10M 0.00025 FroggerNoFrameskip-v4
Pacman Atari Game Image 40M 0.00025 PacmanNoFrameskip-v4
Breakout Atari Game Image 15M 0.00025 BreakoutNoFrameskip-v4
@@
Table 5: Further details for each environment tested in this work. All action spaces were discrete in some form, though for Safety Car a discretized versions of its continuous action space was used. The “Environment Id.” column refers to the environment Id used when generating each environment through the gymnasium interface (Brockman et al., 2016).
Across all our image based domains we utilized a 6x6, checkerboard pattern of 1s and 0s in the top left corner of the image as a trigger. For Safety Car and CAGE-2 we simply append a boolean value to the end of the agent’s observation which we set to 1 in poisoned states, or 0 otherwise. This is also summarized in Table 6. These environments in particular have relatively low dimension and dense state spaces, making the most natural trigger unclear. Thus we implemented this boolean indicator approach for poisoned states to make it it completely clear that there are no collisions between benign and poisoned states in the state space – maintaining our assumption that $S$ and $S_{p}$ are disjoint. Through this design we can better isolate and compare the poisoning strategies of each attack without additional mitigating factors caused by the design of the trigger.
We additionally chose values for $\beta_{low}$ and $\beta_{high}$ to balance attack success and episodic return. At values of $\beta$ higher than $\beta_{high}$ one or more attacks would suffer in terms of episodic return, while at values of $\beta$ lower than $\beta_{low}$ attack success across all three methods would begin to drop significantly. $\beta$ values were chosen on a per-environment basis using the parameters chosen by (Rathbun et al., 2024) as a starting point.
@@ A.2.2 Q-Incept Attack Parameters
Environment Steps per Update Start Poisoning Threshold Trigger
Q*bert 50 6.7% 6X6 Checkerboard
Frogger 50 6.7% 6X6 Checkerboard
Pacman 50 6.7% 6X6 Checkerboard
Breakout 50 6.7% 6X6 Checkerboard
Highway Merge 2 10% 6X6 Checkerboard
Safety Car 4 4.0% Boolean Indicator
CAGE-2 4 4.0% Boolean Indicator
@@
A.3 Further Discussion
@@ A.3.1 Motivation for Baselines
The only other, somewhat comparable attack is BadRL (Cui et al., 2023) which builds upon TrojDRL by optimizing the adversary’s trigger pattern to achieve greater attack success. Trigger optimization is effective but orthogonal to the goals of this work as it can be generically applied to any attack. Furthermore, (Rathbun et al., 2024) showed that BadRL without this trigger fine-tuning often performs worse than TrojDRL, likely since it uses methods to poison the most important – and thus hardest to poison – states in the MDP. Taking all of this into consideration we decided to omit BadRL from our empirical study. Therefore TrojDRL is the best baseline to use when comparing against static reward poisoning attacks using forced action manipulation.
A.3.2 Motivation for Environments
In this paper we study 4 environments – Q*Bert, Frogger, Safety Car, CAGE-2, and Highway Merge. In the TrojDRL paper the authors focused their empirical studies towards Atari game tasks in the gym API (Brockman et al., 2016). We think it is useful to include some of these environments like Q*Bert, as they are standard baselines for RL in general, however we believe it is critical to extend this study to further domains when studying the potential impacts of adversarial attacks. This belief is supported by the findings of (Rathbun et al., 2024) who showed that TrojDRL – which consistently attains near 100% ASR on Atari environments without bounded reward poisoning constraints – often fails to achieve high ASR when tested on non-Atari environments.
@@ A.3.2 Motivation for Environments
A.3.3 Comments on Compute Overhead
Training the Q-network used in Q-Incept does cause some computational overhead which may be detectable, but fortunately it isn’t too extreme. We ran tests on a desktop machine (2x RTX 4090, Threadripper 7980x) and found that SleeperNets, TrojDRL, and Q-Incept run at 1038, 987, and 730 simulation steps per second respectively against Atari Q*bert. However, our Q-network training runs in series with our PPO training, so it’s likely that significant performance increases can be found for Q-Incept by training in parallel.
A.4 Further Experimental Results and Analysis
A.4.1 Q-Incept Against DQN
In Table 7 we present some preliminary results of Q-Incept against DQN. The results show that Q-Incept works well against both on policy (PPO) and off policy (DQN) methods. A better choice of metric $\mathcal{F}_{\hat{Q}}$ may be necessary to maximize performance against DQN, however. For this experiment we slightly alter the methodology of Q-Incept. Instead of poisoning data after each episode and storing it in the agent’s replay buffer, we instead poison data points after they’re sampled from the replay buffer. The reason for this is that the agent’s policy along with the Q-Network of Q-Incept change throughout training, therefore datapoints that were poisoned early in training may become less effective as training progresses. In the case of PPO this wasn’t an issue, since trajectories are only used once and then discarded. For DQN, however, trajectories are stored over a longer time horizon during training, so we need to account for this as an attacker. Equivalently the attacker could keep track of all the data points it has poisoned and properly update them over time.
Beta ASR StDev(ASR) BR StDev(BR)
0% (No Poisoning) N/a N/a 13,724 974
0.5% 92.6% 4.1% 14,092 1,127
1.0% 88.5% 3.4% 13,574 910
Table 7: Performance of Q-Incept against DQN agents training on Atari Q*bert. Results are averaged over 5 runs.
A.4.2 Q-Incept at Different Poisoning Rates
In this section we evaluate the Attack Success Rate performance of Q-Incept at lower $\beta$ values along with the stability of the agent’s Benign Return at higher $\beta$ values. First, in Table 8 we evaluate Q-Incept on Atari Q*bert at $\beta$ values lower than those presented in the main body of the paper. We can see that even at a poisoning rate as low as $\beta=0.03\%$ Q-Incept is still able to achieve near 100% ASR. This is in line with the performance of methods like SleeperNets which presented similar results at $\beta=0.03\%$ while operating under unconstrained rewards. $\beta=0.03\%$ seems to be the minimum necessary poisoning rate for Q-Incept on Q*bert however, as the attack success rate drops off dramatically at $\beta=0.01\%$.
Beta ASR StDev(ASR) BR StDev(BR)
0.3% 100% 0% 18,381 882
0.1% 100% 0% 17,749 1,380
0.05% 100% 0% 17,937 1,304
0.03% 98.3% 2.9% 16,573 873
0.01% 21.1% 6.2% 16,374 2,088
Table 8: Performance of Q-Incept at lower poisoning rates against a PPO agent training on Q*bert.
In Table 9 we compare Q-Incept to the baselines on the CAGE-2 environment to study their stability at higher $\beta$ values. We see that Q-Incept is the only method that improves in ASR as $\beta$ increases, and is also the most stable in terms of BR – never falling below -50.98%. In contrast, SleeperNets and TrojDRL are both highly inconsistent in terms of BR – falling to $-57.82$ and $-64.41$, respectively – while also failing to achieve an ASR above $6\%$, even at $\beta=2\%$.
$\beta$ 0.5% 1% 1.5% 2%
Metric ASR $\sigma$ ASR $\sigma$ ASR $\sigma$ ASR $\sigma$
Q-Incept 30.06% 15.01% 93.21% 15.13% 100% 0.00% 98.62% 2.14%
SleeperNets 0.00% 0.00% 0.06% 0.12% 0.54% 0.89% 1.86% 3.21%
TrojDRL 0.00% 0.00% 5.64% 7.73% 2.24% 3.74% 5.11% 4.99%
Metric BR $\sigma$ Br $\sigma$ BR $\sigma$ BR $\sigma$
Q-Incept -48.64 16.68 -45.49 8.04 -50.98 7.37 -49.41 7.50
SleeperNets -57.82 8.38 -50.18 8.52 -41.44 11.59 -50.67 9.16
TrojDRL -39.28 9.60 -53.11 8.71 -52.77 13.54 -64.41 10.63
Table 9: Comparison of Q-Incept, SleeperNets, and TrojDRL on CAGE at different values of $\beta$.
A.4.3 Q-Incept with an Oracle Q-Network
We noticed that Highway Merge was the only environment on which Q-Incept was unable to attain an average ASR above 90%, leading us to question if our Q-function based approach was incorrect or if our online DQN approximation $\hat{Q}$ wasn’t converging quickly enough. To test this we devised Oracle-Incept – which uses an oracle Q-function pre-trained with DQN until convergence – as a hypothetical, stronger attack by an adversary with direct access to the benign MDP. In Figure 7 we can see that Oracle-Incept improves greatly over Q-Incept, reaching an average ASR of 93.38%. This indicates that better Q-function approximations lead to better performance - validating that both our chosen metric and attack approach scale properly with the accuracy of $\hat{Q}$. Thus, adversaries capable of using Q-function estimations with faster convergence can expect greater attack success.
Figure 7: Comparison between Q-Incept, Oracle-Incept, TrojDRL, and SleeperNets on Highway Merge at $\beta=10\%$. We can see that the Oracle-Incept shows significant improvements in ASR.
A.4.4 Training Plots for Remaining Environments
Here in Figure 8, Figure 9, Figure 10 and Figure 11 we present the training curves for TrojDRL, Q-Incept, and SleeperNets on the Frogger, Pacman, Breakout and Safety Car environments respectively. Plots for Q*Bert are found in Figure 4, CAGE-2 can be found in Figure 3, and Highway Merge in Figure 7. We can see that all attacks perform similarly over time in terms of episodic return on both environments, but Q-Incept is the only attack to reach 100% ASR on average in both environments – doing so very quickly.
Figure 8: Performance of TrojDRL, Q-Incept, and SleeperNets on Frogger at $\beta=0.3\%$
Figure 9: Performance of TrojDRL, Q-Incept, and SleeperNets on Pacman at $\beta=0.3\%$
Figure 10: Performance of TrojDRL, Q-Incept, and SleeperNets on Breakout at $\beta=0.3\%$
Figure 11: Performance of TrojDRL, Q-Incept, and SleeperNets on the Safety Car environment at $\beta=0.1\%$
A.4.5 How Often Does Q-Incept Alter the Agent’s Actions?
Here we explore how often Q-Incept chosen actions in the agent’s replay memory $\mathcal{D}$. We measure this ratio as the number of actions changed divided by the total number of timesteps the attack has poisoned. In Figure 12 and Figure 13 we see that the attack generally balances its action poisoning over time, altering actions on roughly 50% of the time steps it poisons. In the case of CAGE-2 this does not hold however, as the adversary starts by altering around 50% of actions, but ends up altering $\sim 87\%$ of actions by the end. To us this indicates that the difference in values between good and bad actions was much larger in CAGE-2 than in other environments, and furthermore that the agent was highly likely to choose these actions over others as training progressed. Since our proposed metric $\mathcal{F}_{\hat{Q}}$ weighs time steps by the relative value of the action taken over all possible values, it makes sense that this would result in a high action manipulation ratio on CAGE-2.
Figure 12: Ratio of actions changed across poisoned states in our four Atari environments – (from top left to bottom right) Q*bert, Frogger, Pacman, and Breakout. Values are measured as the ratio of actions changed to the total number of poisoned timesteps.
Figure 13: Ratio of actions changed across poisoned states in our three non-Atari environments – (frop top left to bottom) CAGE, Highway Merge, and Safety Car. Values are measured as the ratio of actions changed to the total number of poisoned timesteps.